OpenStudy (dannyo19):

How many different strings of binary length 6 are there such that no two are reversals of each other or add up to 111111?

OpenStudy (noahbred):

69

OpenStudy (dannyo19):

It's not 69.

OpenStudy (noahbred):

why?

OpenStudy (lanhikari22):

Wouldn't 111111 be produced by by only two unique combinations? If so, wouldn't that mean 2^6 - 1, because all combinations produce different strings resulting in not 111111 except for that special case? I'm not sure.

OpenStudy (lanhikari22):

In case of adding two binary numbers, is what I mean.

OpenStudy (zzr0ck3r):

every string has a reverse

OpenStudy (zzr0ck3r):

so 2^5 (without the 111111 thing). Right?

OpenStudy (lanhikari22):

So you divided by two because a whole half of the possibilities is reversed strings?

OpenStudy (zzr0ck3r):

Also, 111111+000000=111111 101010+010101=111111 ....

OpenStudy (zzr0ck3r):

I think....

OpenStudy (lanhikari22):

I don't get the question, is it to be interpreted that we add two binary strings, A and B, and then see the possibility of A + B ending up to be 111111?

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

scroll down for solutions

OpenStudy (lanhikari22):

the number of all possible outputs that can be produced = (2^6)^2 because we can map any possibility with another, much like multiplication, right? the possibility of the two numbers being inverted then is inversionOccurance/totalPossibility

OpenStudy (dannyo19):

How could you prove it's 69?

OpenStudy (dannyo19):

I get 28, but that's not one of the so printed possible answers.

OpenStudy (dannyo19):

this is 2^6 - 8 possible reversable combos all divided by 2 because each can be inverted. Am I right?

OpenStudy (anonymous):

The total number of symmetric \(6\)-digit strings (reversals) is \(2^3=8\) because only the first three digits are needed to determine if a string is symmetric; since this is a manageably small number, I'll list them: \[\begin{matrix}\color{red}{000\,000}&\color{blue}{100\,001}&\color{green}{010\,010}&001\,100\\ 110\,011&\color{green}{101\,101}&\color{blue}{011\,110}&\color{red}{111\,111}\end{matrix}\]Adding the matching-color strings gives \(111\,111\). Meanwhile, there are \(64\) total ways to add binary strings to add up to \(111\,111\) (how many ways can you add two positive integers to add up to \(n\)?). The total number of ways to satisfy \(x\) OR \(y\) is given by the inclusion/exclusion principle: \[n(x\vee y)=n(x)+n(y)-n(x\wedge y)\]